An iron ball diameter measurement is installed automatically in a factory. The gauge will only pass the ball diameter of 1.50 ± d cm. It is known that the ball of the factory production has a diameter that is normally distributed with a mean diameter of 1.50 and a standard deviation of 0.2 cm. If it is desired that 95% of the production passes the selection, should the value be determined?
1
Expert's answer
2020-04-22T18:36:28-0400

let x=diameter of a ball

"X \backsim N(1.5,0.2)"

this can be convert to standard normal by,

"z=\frac{x-\bar{x}}{\sigma}=\frac{x-1.5}{0.2}\\"


here task is to find a find a z value such that,

"P(-z<Z<z)\le0.95\\
2*P(0<Z<z)\le0.95\\
P(0<Z<z)\le0.475"

using table,

"z=1.96"


"z=\frac{x-1.5}{0.2}\\
when\ z=1.96\\
1.96=\frac{x-1.5}{0.2}\\
x_{max}=1.5+0.2*1.96=1.5+0.392\\
when\ z=-1.96\\
-1.96=\frac{x-1.5}{0.2}\\
x_{min}=1.5-0.2*1.96=1.5-0.392\\"


Therefore to pass 95 of ball,diameter should be "1.5\pm 0.392" .

value=0.392